# Trajectory-based models in strong-field physics

### v1

@@ 4 Trajectory-based simulations of frustrated tunneling ionization
The term stabilization in strong-field physics refers to a decrease in the ionization yield when the strength of the laser field exceeds a certain value. This phenomenon attracted a lot of attention in the 1980s and 1990s (see Refs. [1, 2] for reviews). Two different mechanisms underlying the stabilization effect were identified: adiabatic stabilization or stabilization in the Kramers-Henneberger regime and interference stabilization. Adiabatic stabilization corresponds to the situation where the laser frequency exceeds the ionization potential, $\omega>I_{p}$, and the states of the discrete spectrum and states of the continuum are coupled with a single photon energy. The mechanism of interference stabilization is different: it is due to multiple Raman-type transitions from the Rydberg states to the common continuum. In interference stabilization ionization suppression occurs due to the destructive interference of the transition amplitudes between these coherently populated states and the continuum.
The study of Ref. [60] that introduced the term frustrated tunneling ionization (FTI) has renewed interest to the production of neutral atoms in ionization by strong laser fields. In Ref. [60] the yield of neutral excited He atoms produced in ionization of a gaseous target by 30 fs laser pulses with a wavelength of 800 nm and intensities up to $10^{15}$ W/cm2 was experimentally studied. It was found that for relatively low laser intensities the observed yield of the excited neutral atoms is about 20 % of the ion yield. For higher laser intensities this yield decreases to about 10 %. These results were compared with the numerical solution of the TDSE and the outcomes of semiclassical simulations. The semiclassical model of Ref. [60] accounts for both the laser and Coulomb fields. In this model, the electrons with negative energies at the end of the laser pulse were treated as captured into bound states. It was shown that the semiclassical distribution over the effective principal quantum number is in quantitative agreement with the TDSE result.
@@
Formation of highly excited neutral H atoms in the interaction of hydrogen molecules with strong laser pulses is investigated in Ref. [117]. Therefore, the products of this interaction are an excited neutral H atom, a singly charged hydrogen ion, and a liberated electron. A process of this kind may be called frustrated double ionization (FDI), since only one of the two electrons of the H2 molecule eventually escapes [117, 118]. In the trajectory-based approach [117] the full four-body Hamiltonian of the system was propagated in time. It should be stressed that the previously developed 3D semiclassical model did not account for the nuclei motion. Moreover, a transformation to a system of “regularized" coordinates [119] that eliminates the Coulomb singularity was applied in Ref. [117]. The ionization rate of the first electron was described by the semiclassical formula of Ref. [120]. The initial vibrational state of the nuclei is taken to be the ground state of the Morse potential. The Wigner distribution of this state was applied in order to describe the initial state of the nuclei. It was for the first time shown that accounting for two-electron effects is important for the understanding of all pathways leading to the formation of highly excited H atoms [117]. Two such pathways were identified in Ref. [117]: the electron that is released from the molecule either escapes very quickly, or after it remains bound for several periods of the laser pulse, see Fig. 4. It is significant that these two pathways have distinct imprints on the momentum distributions of liberated electrons.
@@
The toolkit for semiclassical simulations of molecules in strong laser fields was developed in Ref. [118]. The computational approach used in [118] for the solution of the system of Newton’s equations of motion is based on the regularization scheme, as well as on the combination of the leapfrog propagation technique with the Bulirsch-Stoer method [121] (see Ref. [122] for a textbook treatment). This method also fully accounts for the Coulomb singularity and is numerically more stable than the one used in [117]. This is due to the fact that the masses do not enter into the time transformation, which leads to a more accurate treatment of many-body systems with large mass ratios. Furthermore, the method of Ref. [118] allows for tunneling of each of the two electrons during the time propagation, i.e., the motion of both electrons is not entirely classical. This feature is important for an accurate description of the enhanced ionization process. Indeed, in this process a double potential well is formed at a critical internuclear distance. As a result, it is easier for an electron bound in the higher potential well to tunnel to the lower potential well and then leave the molecule in the ionization process. The developed toolkit was used to study the FDI process in an elliptically polarized laser field. It is shown that the two-electron effects disappear with increasing ellipticity [118].
[…]
The toolkit [118] turned out to be very powerful in studies of the FDI process. It was applied to the control of electron-electron correlation in FDI with two orthogonally polarized two-color laser fields [124, 125], the study of the FDI process in a heteronuclear HeH+ molecule [126], and enhancing FDI with no electronic correlation using counter-rotating two-color circular laser fields [127]. The study of Ref. [124] focuses on the control of electron-electron correlation in FDI of the D+${}_{3}^{+}$ molecule by using two orthogonally polarized two-color laser fields. It is shown that the control over this process can be achieved by varying the time delay between the two laser pulses: the probability of the FDI process changes sharply with this delay. A characteristic split in the momentum distribution of the liberated electron was identified. The split occurs for time delays corresponding to minima of the FDI probability. It is shown that the split is a signature of the absence of electron-electron correlation in FDI [124]. The sub-cycle control of DFI of the D+${}_{3}^{+}$ molecule was achieved in Ref. [125] using two orthogonally polarized two-color laser fields. In doing so, the FDI process triggered by a strong near-infrared laser field is probed by a weak mid-infrared field. The FDI probability and the momentum distribution of the released electron along the polarization direction of the mid-infrared field considered in conjunction show clear signatures of the sub-cycle control of the process [125]. It is shown that momentum distributions of the escaping electron have a characteristic hive-like shape. The features of this shape can be mapped to the time that one of the two electrons tunnels out of the molecule at the start of the breakup process.
The FDI process in a heteronuclear HeH+ molecule was studied in Ref. [126]. The distributions over the sum of the final nuclei kinetic energies and the principal quantum number were compared to the corresponding distributions for the H2 molecule. It is found that the distribution over the sum of the final kinetic energies of the nuclei has more than one peak for HeH+ interacting with a strong laser pulse. This feature is absent for the H2 molecule. Similarly, the distribution over $n$ has several peaks for HeH+, whereas only a single peak and a “shoulder" are present in this distribution for the H2 molecule. This latter feature is a clear signature of the intertwined electron-nuclear motion. A significant enhancement of the FDI process in the D3+ molecule was demonstrated in Ref. [127] with counter-rotating two-color circular (CRTC) laser fields. A pathway of FDI that is absent for molecules interacting with strong linearly polarized laser field was identified and analyzed [127]. It is found that it is this pathway, in which the first ionization step is frustrated and there is no electron-electron correlation, that is responsible for the enhancement of the process in CRTC fields.
The computational method [118] aimed at description of the H2 and D+${}_{3}^{+}$ molecule was further generalized to more complex atomic and molecular systems in Refs. [128] and [129]. It should be noted that the main problem with 3D semiclassical methods aimed at the description of multielectron ionization and accounting for the Coulomb potential is unphysical autoionization. This process occurs in the following way: one of the bound electrons treated classically can undergo a close encounter with the ionic core and gain a large negative energy. The energy released in this process leads to the release of the other bound electron. Quantum mechanical approaches do not show this problem due to the lower bound of the electron energy. The unphysical autoionization in 3D trajectory-based models can be eliminated by adding a Heisenberg potential [130]. The Heisenberg potential leads to adding a potential barrier that prevents electrons from a close encounter of the core. The potential barrier mimics the Heisenberg uncertainty principle. However, the momentum of a particle is no longer related to the derivative of the position vector in this approach [131, 132]. This is a consequence of adding to the Hamiltonian an extra term depending on position and momentum.
A 3D semiclassical model of triple ionization and frustrated triple ionization of the linear HeH+${}_{2}^{+}$ molecule was developed in Ref. [128]. This model fully accounts for the Coulomb singularities, and it avoids unphysical autoionization by switching off the Coulomb force between two bound electrons. The Coulomb force is then switched on when the motion of one of the electrons is mainly determined by the laser field. The corresponding criteria were formulated [128]. By using this model, both triple and frustrated triple ionization were investigated in Ref. [128]. In the process of frustrated triple ionization (FTI), two electrons are released from an atom or molecule while the third one remains bound, populating a Rydberg state. It is found that similar to the FDI process, two different pathways dominate in FTI. It is also more probable that the electron bound in a Rydberg state is more likely attached to He2+ [128]. It is shown that electronic correlation is weak in both triple ionization and FTI.
The approach [128] was further modified and generalized to the case of three-electron atoms in Ref. [129]. More specifically, the interaction between a pair of bound electrons is described by an effective Coulomb potential [133]. Simultaneously, the Coulomb interaction between the electrons that are not bound is fully taken into account. Therefore, the model [129] is referred to as effective Coulomb potential for bound-bound electrons (ECBB). An efficient set of criteria that are capable of determining whether an electron is bound or “quasifree" during time propagation (i.e., on the fly) was developed in Ref. [129]. The ECBB model can be used for studies of atoms with more than three electrons. It should be mentioned that the model is formulated in the nondipole approximation, i.e., it accounts for the magnetic component of the Lorentz force. The ECBB model was applied to triple and double ionization of the Ar atom by a strong few-cycle laser pulse. It is found that predictions of the model are in good agreement with experimental results for double ionization of Ar [134].
5 Semiclassical models and high-order harmonic generation
The number of works focusing on calculations of HHG spectra using semiclassical models is large but not as large as one might expect, see Refs. [135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148]. To the best of our knowledge, the first paper of such kind dates back to 1990 [135], i.e., before the emergence of the three-step model. In Ref. [135] it was suggested to calculate the harmonic spectrum $I\left(\omega\right)$ as the spectral density of the dipole moment function $z\left(t\right)$:
@@ 5 Semiclassical models and high-order harmonic generation
The single-trajectory analysis of the dipole moment and power spectra generated in the ionization of atomic hydrogen was also carried out in Ref. [136]. At the same time, the HHG spectra calculated by using an ensemble of classical trajectories and Eq.  (19) were presented [136]. As in Ref. [135], the initial conditions were sampled based on the microcanonical distribution. The parameters used in [136] correspond to the tunneling regime: $\omega=0.086$ a.u. and $F=0.11-0.14$ a.u. (the Keldysh parameter $\gamma=0.61-0.78$). The resulting HHG spectra were found to be noisy, and the level of the noise increased with increasing intensity. The origin of this noise was studied [136]. The HHG process in the relativistic regime at the intensities exceeding $3.5\cdot 10^{16}$ W/cm2 was investigated in Ref. [137]. Among other issues, the HHG spectrum in this regime was calculated using a trajectory-based model.
The study of Ref. [138] considers ionization, dissociation, and HHG in the hydrogen molecular ion using a semiclassical model. An interesting approach to the problem of initial conditions for semiclassical simulations based on the Born-Oppenheimer approximation was suggested for the H+${}_{2}^{+}$ ion in Ref. [138]. A more powerful and sophisticated semiclassical model for HHG in hydrogen atom and the H+${}_{2}^{+}$ ion was developed in Ref. [139]. In this model, the electron-nuclei interaction is treated by using a regularization procedure. The procedure relies on transformation to parabolic coordinates and the usage of a fictitious time scale (see Ref. [139] for details). In order to generate the initial microcanonical distribution, action-angle variables were used. HHG spectra calculated using this model demonstrate a characteristic plateau with only odd harmonics. As in Ref. [136], the spectra calculated in [139] become noisier at higher intensities. This was explained by the emergence of chaotic behavior in the atom (ion) exposed to the laser field.
@@
The HHG process in a one-dimensional hydrogen atom was analyzed in Ref. [140]. Such an atom is described by a Hamiltonian $H=p^{2}/2-1/\sqrt{x^{2}+1}+xF\left(t\right)$, where the soft-core parameter $a$ in the potential term is used to remove the Coulomb singularity at $x=0$. A microcanonical distribution [80] was used to sample the initial conditions of classical trajectories. In contrast to Refs. [135, 136, 137, 138, 139], the study [140] deals with the $N$-point discrete representation of the phase-space distribution:
@@
[…]
@@
The study of Ref. [141] focuses on the difference between harmonic spectra calculated using single electron trajectories and the spectra obtained by averaging over an ensemble of classical trajectories. Indeed, the radiation spectrum of a single trajectory contains features that are absent in both experimental observations and fully quantum calculations. On the other hand, if the ensemble averaging is carried out, these unobservable peculiarities disappear, and the resulting high-harmonic spectrum demonstrates all the characteristic HHG features. The latter include the high-energy plateau, cutoff, and the presence of the odd harmonics only. This “ensemble effect" was isolated and thoroughly studied [141]. The study in [141] has found the smallest and simplest subensemble of an ensemble of classical trajectories that is required in order to obtain only the physical features of the HHG process. It should be noted that only relatively low laser intensities corresponding to a Keldysh parameter of about 2.0 are considered in Ref. [141].
@@
Summarizing the above, we note that electrons do not recombine in a classical system, and therefore they do not emit high-order harmonics. This leads to problems in semiclassical simulations of HHG, since the quantum recombination process must be artificially inserted in a trajectory-based model. In modern semiclassical simulations, this is done by using two different approaches. It is either assumed that the returning electron emits a photon with energy equal to the sum of the instantaneous energy of the electron and the ionization potential [145, 146, 147, 148] or a quantum recombination dipole matrix element is combined with classical density of returning electrons (CTMC-QUEST model). However, both these hybrid approaches introduce some limitations. For example, this applies to producing accurate interference patterns in HHG spectra, which is difficult to achieve using semiclassical models due to the lack of the phase information. Surprisingly, there is a lack of studies comparing the semiclassical simulations with the TDSE results for the generation of high-order harmonics [149]. Therefore, further developments of trajectory-based models for HHG are needed.
@@ 6.2 Nonsequential double ionization of atoms by elliptically polarized fields
For the first time, the ion and electron momentum distributions in NSDI of a model atom by an elliptically polarized laser field were calculated in Ref. [55]. To this end, a simple semiclassical model based on tunneling, electron kinematics in the laser field only, and phase space volume was used in [55]. Electron-electron interaction was modelled as a three-body contact interaction located at the position of the ion (see Ref. [55] for details). It was found that for ellipticities exceeding approximately $0.3$, the shortest electron trajectory, i.e., the trajectory that has the shortest travel time before return to an ion, no longer dominates the double ionization rate, which is in contrast to the case of linear polarization. Simultaneously, significant violations of the symmetry patterns in the ion-momentum distributions and the electron momentum correlations were found, compared with the case of a linearly polarized field, see Fig. 7. Since the late returns are absent for very short pulses, these violations are sensitive to the laser pulse duration. It was pointed out that the predicted qualitative effects in electron distributions should be experimentally detectable, and their observation would also verify the significance of the late returns of the ionized electron to the parent ion [55]. To date, a plethora of works exploring various aspects of NSDI in elliptically and circularly polarized fields has appeared, see e.g., Refs. [57, 56, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176]. All these studies are based on some semiclassical models.
@@
In Ref. [56], the NSDI process in an elliptically polarized field was investigated by using the model accounting for the electron motion in both laser and soft-core Coulomb fields. The electron-electron interaction was also simulated with the soft-core Coulomb potential. While the study [55] was restricted to ellipticity smaller than $0.5$, momentum distributions of doubly charged ions were studied for ellipticities from 00 to $1$ in Ref. [56]. Therefore, the evolution of the ion momentum distributions when polarization changes from linear to circular was studied for the first time [56]. It was shown that the contributions of sequential and nonsequential double ionization in the ion momentum distributions can be clearly distinguished. It was found that the ellipticity of the strong laser pulse is closely related to the timing of the NSDI process [167]. More specifically, the analysis of classical trajectories has shown that higher ellipticity forces the beginning of NSDI into a narrower time window. In turn, this “pins" the field strength of the ionizing field in an unexpected way [167]. By using a trajectory-based model, it was shown that the NSDI process can exist even in circularly polarized laser fields [168]. It turns out that the underlying mechanism is again based on recollision. NSDI in a circularly polarized laser was also studied in Ref. [169] for Mg atoms. Semiclassical simulations employing the soft-core Coulomb potential have shown that for circular polarization only a single recollision event leads to the NSDI process.

### v2

@@ 4 Trajectory-based simulations of frustrated tunneling ionization
In strong-field physics, stabilization describes the reduction of ionization yield at laser field strengths above a certain threshold. This phenomenon attracted a lot of attention in the 1980s and 1990s (see Refs. [1, 2] for reviews). Stabilization has been attributed to two mechanisms: adiabatic (Kramers-Henneberger) stabilization and interference stabilization. Adiabatic stabilization corresponds to the situation where the laser frequency exceeds the ionization potential, $\omega>I_{p}$, and the states of the discrete spectrum and states of the continuum are coupled with a single photon energy. The mechanism of interference stabilization is different: it is due to multiple Raman-type transitions from the Rydberg states to the common continuum. In interference stabilization, ionization is suppressed through destructive interference between transition amplitudes connecting coherently populated Rydberg states to the continuum.
The study of Ref. [60] that introduced the term frustrated tunneling ionization (FTI) stimulated further investigations of the production of neutral atoms in ionization by strong laser fields. In Ref. [60] the yield of neutral excited He atoms produced in ionization of a gaseous target by 30 fs laser pulses with a wavelength of 800 nm and intensities up to $10^{15}$ W/cm2 was experimentally studied. For moderate laser intensities, the yield of excited neutral atoms was around 20% of the ion yield, decreasing to  10% at higher intensities. These findings were compared with TDSE calculations and semiclassical simulations. The semiclassical model [60] includes both laser and Coulomb interactions, and electrons remaining with negative energies after the pulse are considered captured into bound states. It was shown that the semiclassical distribution over the effective principal quantum number is in quantitative agreement with the TDSE result.
@@
Formation of highly excited neutral H atoms in the interaction of hydrogen molecules with strong laser pulses is investigated in Ref. [117]. Therefore, the products of this interaction are an excited neutral H atom, a singly charged hydrogen ion, and a liberated electron. A process of this kind may be called frustrated double ionization (FDI), since only one of the two electrons of the H2 molecule eventually escapes [117, 118]. In Ref. [117], the full four-body Hamiltonian was propagated in time using a regularization of the coordinates [119] to eliminate Coulomb singularities. Importantly, the previously developed 3D semiclassical model did not account for the nuclei motion. The ionization rate of the first electron was described by the semiclassical formula of Ref. [120]. The initial vibrational state of the nuclei is taken to be the ground state of the Morse potential. The Wigner distribution of this state was applied in order to describe the initial state of the nuclei. These results demonstrate for the first time that accounting for two-electron effects is important for the understanding of all pathways leading to the formation of highly excited H atoms [117]. Two such pathways were identified in Ref. [117]: the electron that is released from the molecule either escapes very quickly, or after it remains bound for several periods of the laser pulse, see Fig. 4. Notably, these two pathways have distinct imprints on the momentum distributions of liberated electrons.
@@
A computational toolkit enabling semiclassical simulations of molecules in strong laser fields was developed in Ref.  [118]. This method combines leapfrog propagation with the Bulirsch-Stoer scheme [121] (see Ref. [122] for a textbook treatment), regularizes Coulomb singularities, and provides a stable treatment for systems with large mass ratios. Furthermore, the method of Ref. [118] allows for tunneling of each of the two electrons during the time propagation, i.e., the motion of both electrons is not entirely classical. This feature is important for an accurate description of the enhanced ionization process. Indeed, in this process a double potential well is formed at a critical internuclear distance. As a result, it is easier for an electron bound in the higher potential well to tunnel to the lower potential well and then leave the molecule in the ionization process. The developed toolkit was used to study the FDI process in an elliptically polarized laser field. It is shown that the two-electron effects disappear with increasing ellipticity [118].
@@
The FDI process in a heteronuclear HeH+ molecule was studied in Ref. [126]. The distributions over the sum of the final nuclei kinetic energies and the principal quantum number were compared to the corresponding distributions for the H2 molecule. It is found that the distribution over the sum of the final kinetic energies of the nuclei has more than one peak for HeH+ interacting with a strong laser pulse. This feature is absent for the H2 molecule. Similarly, the distribution over $n$ has several peaks for HeH+, whereas only a single peak and a “shoulder” are present in this distribution for the H2 molecule. This latter feature is a clear signature of the intertwined electron-nuclear motion. A significant enhancement of the FDI process in the D3+ molecule was demonstrated in Ref. [127] with counter-rotating two-color circular (CRTC) laser fields. A pathway of FDI that is absent for molecules interacting with strong linearly polarized laser field was identified and analyzed [127]. It is found that it is this pathway, in which the first ionization step is frustrated and there is no electron-electron correlation, that is responsible for the enhancement of the process in CRTC fields.
[…]
The approach [128] was further modified and generalized to the case of three-electron atoms in Ref. [129]. More specifically, the interaction between a pair of bound electrons is described by an effective Coulomb potential [133]. Simultaneously, the Coulomb interaction between the electrons that are not bound is fully taken into account. Therefore, the model [129] is referred to as effective Coulomb potential for bound-bound electrons (ECBB). An efficient set of criteria that are capable of determining whether an electron is bound or “quasifree” during time propagation (i.e., on the fly) was developed in Ref. [129]. The ECBB model can be used for studies of atoms with more than three electrons. The model is formulated within the nondipole approximation, incorporating the magnetic component of the Lorentz force. The ECBB model was applied to triple and double ionization of the Ar atom by a strong few-cycle laser pulse. It is found that predictions of the model are in good agreement with experimental results for double ionization of Ar [134].
5 Semiclassical models and high-order harmonic generation
Although many studies have investigated HHG spectra using semiclassical models, the number is smaller than might be anticipated, see Refs. [135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148]. The earliest study of this type was published in 1990 [135], prior to the development of the three-step model. In Ref. [135] it was suggested to compute the harmonic spectrum $I\left(\omega\right)$ as the spectral density of the dipole moment $z\left(t\right)$:
$I\left(\omega\right)=\frac{1}{2\pi}\lim_{T\rightarrow 0}\frac{1}{2T}\left<\left|\int_{0}^{2\pi}z\left(t\right)\exp\left(i\omega t\right)dt\right|^{2}\right$>, (19)
[…]
@@ 5 Semiclassical models and high-order harmonic generation
The single-trajectory analysis of the dipole moment and power spectra generated in the ionization of atomic hydrogen was also carried out in Ref. [136]. At the same time, the HHG spectra calculated by using an ensemble of classical trajectories and Eq.  (19) were presented [136]. As in Ref. [135], the initial conditions were sampled based on the microcanonical distribution. The parameters used in [136] correspond to the tunneling regime: $\omega=0.086$ a.u. and F=0.110.14F=0.11-0.14 a.u. (the Keldysh parameter $\gamma=0.61-0.78$). The resulting HHG spectra were found to be noisy, and the level of the noise increased with increasing intensity. The origin of this noise was studied [136]. The HHG process in the relativistic regime at the intensities exceeding $3.5\cdot 10^{16}$ W/cm2 was investigated in Ref. [137]. Among other issues, the HHG spectrum in this regime was calculated using a trajectory-based model.
[…]
@@
The HHG process in a one-dimensional hydrogen atom was analyzed in Ref. [140]. Such an atom is described by a Hamiltonian $H=p^{2}/2-1/\sqrt{x^{2}+1}+xF\left(t\right)$, where the soft-core parameter $a$ in the potential term is used to remove the Coulomb singularity at $x=0$. A microcanonical distribution [80] was used to sample the initial conditions of classical trajectories. Unlike earlier studies, Ref. [140] applied a discrete representation of the phase-space distribution for semiclassical simulations:
@@
The density of Eq. (21) is used to calculate the dipole $d_{C}\left(t\right)=\int\rho_{C}\left(t\right)xdx$. The HHG spectrum is obtained as a Fourier transform of the dipole acceleration:
$I\left(\omega\right)=\left|\frac{1}{\sqrt{2\pi}}\int_{0}^{T}\ddot{d}_{C}\left(t\right)\exp\left(-i\omega t\right)dt\right|^{2}$ (22)
@@
The study of Ref. [141] focuses on the difference between harmonic spectra calculated using single electron trajectories and the spectra obtained by averaging over an ensemble of classical trajectories. Single-trajectory spectra display features that are not observed in experiments or quantum calculations; averaging over an ensemble removes these unphysical artifacts, recovering the characteristic HHG features, including the plateau, cutoff, and the presense of the odd harmonics only. This “ensemble effect” was isolated and thoroughly studied [141]. The study in [141] has found the smallest and simplest subensemble of an ensemble of classical trajectories that is required in order to obtain only the physical features of the HHG process. It should be noted that only relatively low laser intensities corresponding to a Keldysh parameter of about 2.0 are considered in Ref. [141].
@@
Summarizing the above, we note that electrons do not recombine in a classical system, and therefore they do not emit high-order harmonics. This leads to problems in semiclassical simulations of HHG, since the quantum recombination process must be artificially inserted in a trajectory-based model. In modern semiclassical simulations, this is done by using two different approaches. It is either assumed that the returning electron emits a photon with energy equal to the sum of the instantaneous energy of the electron and the ionization potential [145, 146, 147, 148] or a quantum recombination dipole matrix element is combined with classical density of returning electrons (CTMC-QUEST model). However, both these hybrid approaches introduce some limitations. For example, this applies to producing accurate interference patterns in HHG spectra, which is difficult to achieve using semiclassical models due to the lack of the phase information. There is a lack of studies comparing the semiclassical simulations with the TDSE results for the generation of high-order harmonics [149]. Therefore, further developments of trajectory-based models for HHG are needed.
@@ 6.2 Nonsequential double ionization of atoms by elliptically polarized fields
The first calculations of ion and electron momentum distributions in NSDI of a model atom driven by an elliptically polarized laser field were presented in Ref. [55]. In that study, a simple semiclassical framework was employed, combining tunneling ionization, classical propagation of the electron in the laser field, and phase-space considerations. The electron–electron interaction was approximated by a localized three-body contact potential at the ionic core (see details in Ref. [55]). It was demonstrated that when the ellipticity exceeds about 0.3, the contribution of the shortest electron trajectory—corresponding to the fastest return to the ion—ceases to dominate the double-ionization yield, in contrast to the situation under linear polarization. This change is accompanied by pronounced asymmetries in both ion momentum distributions and electron–electron correlations, as illustrated in Fig. 7. Since late electron returns are suppressed in ultrashort pulses, these asymmetries were found to depend sensitively on the pulse duration. The study also suggested that the predicted features in the electron momentum distributions should be observable experimentally, thus providing evidence for the role of late electron returns. To date, a plethora of works exploring various aspects of NSDI in elliptically and circularly polarized fields has appeared, see e.g., Refs. [57, 56, 167, 168, 169, 170, 171, 172, 173, 174, 175, 176]. All these studies are based on some semiclassical models.
@@
In Ref. [56], the NSDI process in an elliptically polarized field was investigated by using the model accounting for the electron motion in both laser and soft-core Coulomb fields. The electron-electron interaction was also simulated with the soft-core Coulomb potential. While the study [55] was restricted to ellipticity smaller than $0.5$, momentum distributions of doubly charged ions were studied for ellipticities from 0 to $1$ in Ref. [56]. Therefore, the evolution of the ion momentum distributions when polarization changes from linear to circular was studied for the first time [56]. It was shown that the contributions of sequential and nonsequential double ionization in the ion momentum distributions can be clearly distinguished. It was found that the ellipticity of the strong laser pulse is closely related to the timing of the NSDI process [167]. More specifically, the analysis of classical trajectories has shown that higher ellipticity forces the beginning of NSDI into a narrower time window. In turn, this “pins” the field strength of the ionizing field in an unexpected way [167]. By using a trajectory-based model, it was shown that the NSDI process can exist even in circularly polarized laser fields [168]. It turns out that the underlying mechanism is again based on recollision. NSDI in a circularly polarized laser was also studied in Ref. [169] for Mg atoms. Semiclassical simulations employing the soft-core Coulomb potential have shown that for circular polarization only a single recollision event leads to the NSDI process.